Using the principle of induction, prove that 64 is a factor of 3^(2n+2)- 8n-9 ∀ n∈N

1
Expert's answer
2022-01-28T13:37:51-0500

Let "P(n)" be the proposition that "64" is a factor of "3^{2n+2}-8n-9, \forall n\in \N."

Basis Step

"P(1)" is true because "64" is a factor of "3^{2(1)+2}-8(1)-9=64."

Inductive Step

We assume that "64" is a factor of "3^{2k+2}-8k-9, k\in \N."

Under this assumption, it must be shown that "P(k + 1)" is true, namely, that

"64" is a factor of "3^{2(k+1)+2}-8(k+1)-9, k\in \N."


"3^{2(k+1)+2}-8(k+1)-9=9\cdot3^{2k+2}-8k-17"

"=9\cdot3^{2k+2}-72k+64k-81+64"

"=9(3^{2k+2}-8k-9)+64(k+1)"

This last equation shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for every natural "n". That is, we have proved that "64" is a factor of "3^{2n+2}-8n-9, \forall n\in \N."